# Empowering 3D Visual Grounding with Reasoning Capabilities

### v2

@@ Abstract
Although great progress has been made in 3D visual grounding, current models still rely on explicit textual descriptions for grounding and lack the ability to reason human intentions from implicit instructions. We propose a new task called 3D reasoning grounding and introduce a new benchmark ScanReason which provides over 10K question-answer-location pairs from five reasoning types that require the synerization of reasoning and grounding. We further design our approach, ReGround3D, composed of the visual-centric reasoning module empowered by Multi-modal Large Language Model (MLLM) and the 3D grounding module to obtain accurate object locations by looking back to the enhanced geometry and fine-grained details from the 3D scenes. A chain-of-grounding mechanism is proposed to further boost the performance with interleaved reasoning and grounding steps during inference. Extensive experiments on the proposed benchmark validate the effectiveness of our proposed approach. Codes, datasets, and benchmarks will be available at https://zcmax.github.io/projects/ScanReason.
@@ 1 Introduction
Figure 1: For an embodied agent, they not only need to be able to understand the 3D environment and complex human instructions, but also localize the target objects for interaction and navigation. Although GPT-4 (GPT-4V) have strong text (multi-modal) reasoning abilities, they lack the ability to directly perceive the 3D scene, understand the 3D spatial relationships and output corresponding target object locations. Instead, our proposed method ReGround3D has the 3D perception, reasoning, and grounding capabilities in the real-world 3D environment.
@@
Understanding and reasoning in the 3D visual world is critical for applications such as robotics and augmented reality, where embodied agents are expected to understand the 3D layout and predict the 3D locations of objects based on human instructions. The example in Fig. 1 demonstrates a scenario where the question can only be solved with a comprehensive understanding of the 3D scene and joint reasoning of the question and the 3D environment. However, current 3D visual grounding models [23, 38, 49, 42] trained on [6, 1] localize objects based on explicit descriptions of the object category, attribute, and 3D spatial relationships, and lack the ability to reason the user intentions and predict object locations with implicit human instructions such as “I am thirsty, can I have something to drink?”.
@@
To bridge the aforementioned gap and to push the boundaries of what embodied agents can understand and how they can interact with the 3D world, we propose a new task of 3D reasoning grounding and introduce a new benchmark named ScanReason. The task requires the model to conduct joint reasoning on the question and the 3D environment before predicting the 3D locations of target objects. We define five categories of 3D reasoning: spatial reasoning, function reasoning, logistic reasoning, emotional reasoning, and safety reasoning. The first two categories focus on the fundamental understanding of the 3D physical world, while the last three categories are built upon fundamental abilities to address user-centric real-world challenges. The benchmark comprises more than 10K question-answer-3D bounding box pairs from 2K scenes belonging to the five reasoning types mentioned above. The GPT-4-assisted data annotation process largely increases the efficiency of curating such a dataset.
@@
We propose ReGround3D as an initial attempt to the new task of 3D Reasoning Grounding. Intuitively, for 3D grounding with implicit instructions, we need to first conduct reasoning on the language instructions and the coarse visual environment. Then with the idea of which object we want to find in mind, we look back to the 3D scene and ground the target object. For complex instructions we may need to alternate the reasoning and look-back process for multiple iterations. Inspired by this intuition, our framework is composed of visual-centric reasoning module and 3D grounding with geometry-enhanced look-back module, with a Chain-of-Grounding mechanism during inference to alternately conduct reasoning and grounding for multiple for multiple rounds. Specifically, the visual-centric reasoning module conducts joint reasoning of 3D scene and instructions with an MLLM. This module predicts a special token representing the semantic and location information of the target object, which is used for the grounding module. The 3D grounding module uses the output token embedding from the previous reasoning module to locate the target object by looking back to the fine-grained 3D scene representation. Unlike previous MLLMs attempting to directly predict bounding box coordinates, our look-back mechanism enables the model to capture more comprehensive 3D geometry and fine-grained object details for accurate 3D grounding. The Chain-of-Grounding mechanism is proposed to synergize reasoning and grounding, which allows multiple rounds alternating between reasoning and grounding during inference.
@@
In summary, our contributions are threefold: 1) We propose a new task of 3D reasoning grounding which requires the model synergize reasoning and grounding. We further introduce a new benchmark ScanReason, comprises five reasoning types (spatial reasoning, function reasoning, logistic reasoning, emotional reasoning, and safety reasoning) for the task of 3D reasoning grounding in diverse scenes. 2) We design a new framework ReGround3D with a visual-centric reasoning module and a 3D grounding module with geometry-enhanced look-back. We further introduce a Chain-of-Grounding mechanism to boost the 3D reasoning grounding ability with a chain of interleaved reasoning and grounding steps. 3) Extensive experiments demonstrate the effectiveness of the our ReGround3D on the ScanReason benchmark for 3D reasoning grounding.
@@ 2 Related Work
The task of 3D visual grounding is aimed at localizing the objects that are explicitly referred to by free-form guided language expressions in the 3D scene. Inspired by the success of transformers in natural language processing, recent 3D visual grounding approaches [38, 23, 49, 52, 11, 16] have started to adopt transformer [34] architectures for handling the complex relationships between language descriptions and 3D visual data. These methods leverage the self-attention mechanism of transformers to dynamically weigh the importance of different parts of the input data, facilitating more effective grounding of textual descriptions in 3D environment. Compared with 3D visual grounding, our proposed 3D reasoning grounding requires model to reason the complex question, ground target objects and give the explanation at the same time.
@@ 3 ScanReason Benchmark
Given a 3D scene and an implicit free-form query, 3D reasoning grounding requires the model to predict the 3D bounding boxes of the target objects as well as the textual answers and explanations. As shown in Fig. 2, different from traditional 3D visual grounding, the queries of 3D reasoning grounding are implicit and complex, requiring strong reasoning, commonsense, and world knowledge. The number of target objects in 3D reasoning grounding is flexible, and any object in the 3D scene that satisfies the query requirements should be considered as the target object.
Figure 2: The left side figure shows the overall of our ScanReason dataset. For each reasoning category, we designed different prompts to generate corresponding questions. And the right side figure shows the differences between traditional 3D visual grounding task and our proposed 3D reasoning grounding task.
@@ 3.2 Question Types
To comprehensively evaluate the 3D reasoning grounding abilities, we define 5 types of questions depending on which type of reasoning is required. Spatial reasoning and function reasoning require fundamental understanding of the 3D physical world, while logistic reasoning, emotional reasoning, and safety reasoning are high-level reasoning skills built upon the two fundamental reasoning abilities to address user-centric real-world applications, as shown in Fig. 2.
3.2.1 Spatial Reasoning
measures models’ understanding of 3D spatial relationships among objects in a 3D scene. It encompasses the ability to comprehend the layout and structure of the 3D scene, the 3D location of objects within it, which could serve as the foundation of navigating or planning movements in the 3D environment.
3.2.2 Function Reasoning
involves understanding and inferring the purpose, function, or affordance of objects within the 3D scene. For example, function reasoning allows an embodied agent to recognize that a chair is for sitting, a lamp is for lighting, and a refrigerator is for storing food at low temperatures. Such understanding enables the embodied agent to assist users and to perform complex tasks more effectively (e.g., turning on a lamp when the room gets dark, or navigating to a refrigerator to fetch a drink).
3.2.3 Logistic Reasoning
allows an embodied agent to not only understand its environment but also to interact with it in a goal-directed manner. For example, given a question shown in Fig. 2: “If I’m cooking dinner in the kitchen, where is the nearest place for me to throw the rubbish?”, an agent need to use such reasoning ability to infer the location of objects (in this question, a rubbish bin) based on their function and spatial relationships under the specific setting (the kitchen).
3.2.4 Emotional Reasoning
plays a critical role in human-robot interaction, where the target objects are determined by understanding human emotions, preferences, and behavioral patterns. This ability makes the embodied agents more attuned to the emotional and psychological states of humans, allowing them to provide more personalized, empathetic, and contextually appropriate responses and solutions, such as: “I’m very sad, is there something can make me feel happy?” shown in 2.
3.2.5 Safety Reasoning
focuses on preventing harm and ensuring the well-being of humans and other creatures in the 3D environment. It requires the embodied agent to identify and assess the risk and make safety-aware decisions, such as: “Where should first aid kits be placed for easy access but out of reach of children?”.
@@ 3.3 Automatic Data Annotation with GPT-4
In total, our ScanReason dataset consists of 12929 complex reasoning question-answer-location pairs from 1456 scenes, which are split into 11455 training pairs and 1474 validation pairs. All 1474 validation question-answer pairs have been manually verified, including 342 spatial reasoning questions, 287 function reasoning questions, 581 logistic reasoning questions, 211 emotional questions and 53 safety reasoning questions. We provide the detailed statistics and more examples of our dataset in Appendix.
@@ 3.4 Evaluation Metric
To evaluate the accuracy of predicted objects and their locations for a flexible number of ground-truth objects, we follow the evaluation metric of 3D object detection tasks. Specifically, we adopt Acc@$k$IoU as our metric, where $k$ is the threshold for the Intersection of Union (IoU) between positive predictions and ground-truths. We evaluate the performance under $k=0.25$ in our experiments.
@@ 4 Method
Solving the task of 3D reasoning grounding requires the synergization of the perception, reasoning, and grounding capability of the embodied agent. Intuitively, we can first conduct reasoning based on the implicit instruction such as “where should first aid kits be placed?” and the visual environment. The reasoning process provides us with information about the rough location and semantics of the object we are looking for. Then keeping that information in mind, we look back to the 3D environment to precisely locate the object. For complex scenarios, alternate reasoning and looking back are required for multiple rounds until we obtain the final answer.
Inspired by this intuition, we propose ReGround3D, consisting of a visual-centric reasoning module and a 3D grounding module with geometry-enhanced look-back, as illustrated in Fig. 3. The visual-centric reasoning module (Sec. 4.1) performs joint reasoning of language instruction and visual scene, and predicts a special <LOC> token representing the grounding information. The 3D grounding module (Sec. 4.2) looks back to the original 3D scene with comprehensive geometry information and fine-grained details. It takes the hidden embedding of the <LOC> token containing grounding-related information from the 3D features, and eventually predicts the 3D locations of the target objects. Furthermore, we propose Chain-of-Grounding mechanism (CoG) (Sec. 4.3), i.e., a chain of interleaved reasoning and grounding steps, to further synergize the grounding and reasoning capability for the 3D reasoning grounding task, as illustrated in Fig. 4.
@@ 4.1 Visual-Centric Reasoning
Specifically, for simplicity, we leverage 3D-LLM to serve as the visual-centric reasoning module because of its strong reasoning abilities inherited from the LLM. Based on the BLIP2 architecture [25], 3D-LLM uses pre-trained image encoders to extract multi-view 2D image features and unprojects them into 3D spaces. The visual features are encoded by the Q-Former to produce 32 tokens as the visual input to the LLM. By leveraging the pretrained image encoders and the Q-Former, the visual tokens encode rich semantics but lack the 3D structures, spatial interactions, and fine-grained details. Therefore, instead of directly predicting the object locations by the 3D-LLM, we ask the 3D-LLM to predict the feature representation as the output of the reasoning process, and the predicted feature is further used to ground the target object in the grounding module.
@@
4.2.1 3D Visual Encoder
Unlike the 2D image encoder used for 3D-LLM, the 3D visual encoder directly extracts features from 3D point clouds to capture more geometric and spatial information about the 3D structures and fine-grained details. The powerful 3D visual encoder which captures comprehensive geometry, structure, layout, and fine-grained information is critical to accurate 3D grounding. Subsequently, the 3D features $f_{scene}$ produced by the 3D visual encoder and the grounding feature $h_{loc}$ from 3D-LLM are sent to the query selection module.
4.2.2 Query Selection Module
We adopt a cross-attention mechanism, where we treat $f_{scene}$ as $Q$ (Query), and $h_{loc}$ as both $K$ (Key) and $V$ (Value), to implicitly obtain a feature-level reasoning activation heatmap. During this scene look-back process, this module roughly locates scene features that have a high response to the <LOC> token. We then select the $k$ most relevant features as the object query $f_{query}$.
4.2.3 3D Box Decoder
is a classical transformer decoder, which consists of M transformer decoder layers, In each decoder layer, the object queries $f_{query}$ go through the text feature $h_{loc}$ cross-attention layer and scene feature $f_{scene}$ cross-attention layer. Finally, the prediction head takes the updated object queries as input and predicts the final 3D locations and matching score.
4.2.4 Discussion
In comparison to previous works [10, 14, 20, 12] that directly predicts the bounding boxes by the LLM, the extra grounding module has the following advantages: 1) Based on the MLLM reasoning results, it allows the grounding module to perceive the scene again and focus on the region under the implicit guidance of MLLM, adapted to the user queries and the reasoning results. 2) The scene representation perceived in the grounding module can be more precise and fine-grained, which is complementary to the visual-centric reasoning module. 3) The two-step reasoning-grounding pipeline is flexible and can generalized to other types of predicting formats such as segmentation masks (by simply replacing the 3D box decoder with a 3D mask decoder).

### v3

# https://zcmax.github.io/projects/ScanReason

@@ Abstract
Although great progress has been made in 3D visual grounding, current models still rely on explicit textual descriptions for grounding and lack the ability to reason human intentions from implicit instructions. We propose a new task called 3D reasoning grounding and introduce a new benchmark ScanReason which provides over 10K question-answer-location pairs from five reasoning types that require the synerization of reasoning and grounding. We further design our approach, ReGround3D, composed of the visual-centric reasoning module empowered by Multi-modal Large Language Model (MLLM) and the 3D grounding module to obtain accurate object locations by looking back to the enhanced geometry and fine-grained details from the 3D scenes. A chain-of-grounding mechanism is proposed to further boost the performance with interleaved reasoning and grounding steps during inference. Extensive experiments on the proposed benchmark validate the effectiveness of our proposed approach.
@@ 1 Introduction
Figure 1: For an embodied agent, they not only need to be able to understand the 3D environment and complex human instructions but also localize the target objects for interaction and navigation. Although GPT-4 (GPT-4V) have strong text (multi-modal) reasoning abilities, they lack the ability to directly perceive the 3D scene, understand the 3D spatial relationships, and output corresponding target object locations. Instead, our proposed method ReGround3D has the 3D perception, reasoning, and grounding capabilities in the real-world 3D environment.
@@
Understanding and reasoning in the 3D visual world is critical for applications such as robotics and AR, where embodied agents are expected to understand the 3D layout and predict the 3D locations of objects based on human instructions. The example in Fig. 1 demonstrates a scenario where the question can only be solved with a comprehensive understanding of the 3D scene and joint reasoning of the question and the 3D environment. However, current 3D visual grounding models [24, 39, 50, 43] trained on [7, 1] localize objects based on explicit descriptions of the object category, attribute, and 3D spatial relationships, and lack the ability to reason the user intentions and predict object locations with implicit human instructions such as “I am thirsty, can I have something to drink?”.
@@
To bridge the aforementioned gap and to push the boundaries of what embodied agents can understand and how they can interact with the 3D world, we propose a new task of 3D reasoning grounding and introduce a new benchmark named ScanReason. The task requires the model to conduct joint reasoning on the question and the 3D environment before predicting the 3D locations of target objects. We define five categories of 3D reasoning: spatial reasoning, functional reasoning, logical reasoning, emotional reasoning, and safety reasoning. The first two categories focus on the fundamental understanding of the 3D physical world, while the last three categories are built upon fundamental abilities to address user-centric real-world challenges. The benchmark comprises more than 10K question-answer-3D bounding box pairs from 2K scenes belonging to the five reasoning types mentioned above. The GPT-4-assisted data annotation process largely increases the efficiency of curating such a dataset.
@@
We propose ReGround3D as an initial attempt to the new task of 3D Reasoning Grounding. Intuitively, for 3D grounding with implicit instructions, we need to first conduct reasoning on the language instructions and the coarse visual environment. Then with the idea of which object we want to find in mind, we look back to the 3D scene and ground the target object. For complex instructions, we may need to alternate the reasoning and look-back process for multiple iterations. Inspired by this intuition, our framework is composed of a visual-centric reasoning module and 3D grounding with geometry-enhanced look-back module, with a Chain-of-Grounding mechanism during inference to alternately conduct reasoning and grounding for multiple rounds. Specifically, the visual-centric reasoning module conducts joint reasoning of the 3D scene and instructions with an MLLM. This module predicts a special token representing the semantic and location information of the target object, which is used for the grounding module. The 3D grounding module uses the output token embedding from the previous reasoning module to locate the target object by looking back at the fine-grained 3D scene representation. Unlike previous MLLMs attempting to directly predict bounding box coordinates, our look-back mechanism enables the model to capture more comprehensive 3D geometry and fine-grained object details for accurate 3D grounding. The Chain-of-Grounding mechanism is proposed to synergize reasoning and grounding, which allows multiple rounds alternating between reasoning and grounding during inference.
@@
In summary, our contributions are threefold: 1) We propose a new task of 3D reasoning grounding which requires the model to synergize reasoning and grounding. We further introduce a new benchmark ScanReason, which comprises five reasoning types (spatial reasoning, functional reasoning, logical reasoning, emotional reasoning, and safety reasoning) for the task of 3D reasoning grounding in 3D scenes. 2) We design a new framework ReGround3D with a visual-centric reasoning module and a 3D grounding module with geometry-enhanced look-back. We further introduce a Chain-of-Grounding mechanism to boost the 3D reasoning grounding ability with a chain of interleaved reasoning and grounding steps. 3) Extensive experiments demonstrate the effectiveness of the our ReGround3D on the ScanReason benchmark for 3D reasoning grounding.
@@ 2 Related Work
The task of 3D visual grounding is aimed at localizing the objects that are explicitly referred to by free-form guided language expressions in the 3D scene. Inspired by the success of transformers in natural language processing, recent 3D visual grounding approaches [39, 24, 50, 53, 12, 17] have started to adopt transformer [35] architectures for handling the complex relationships between language descriptions and 3D visual data. These methods leverage the self-attention mechanism of transformers to dynamically weigh the importance of different parts of the input data, facilitating a more effective grounding of textual descriptions in the 3D environment. Recent method [3] proposes a Chain-of-Thoughts module that predicts a chain of anchor objects that are subsequently utilized to localize the final target object. Compared with 3D visual grounding, our proposed 3D reasoning grounding requires the model to reason the complex question, ground target objects, and give the explanation at the same time.
@@ 3 ScanReason Benchmark
Given a 3D scene and an implicit free-form query, 3D reasoning grounding requires the model to predict the 3D bounding boxes of the target objects as well as the textual answers and explanations. As shown in Fig. 2, different from traditional 3D visual grounding, the queries of 3D reasoning grounding are implicit and complex, requiring strong reasoning, commonsense, and world knowledge. The number of target objects in 3D reasoning grounding is flexible and any object satisfying the requirements should be considered as the target object.
Figure 2: The left side figure shows the overall of our ScanReason dataset. For each reasoning category, we designed different prompts to generate corresponding questions. And the right side figure shows the differences between the traditional 3D visual grounding task and our proposed 3D reasoning grounding task.
@@ 3.2 Question Types
To comprehensively evaluate the 3D reasoning grounding abilities, we define 5 types of questions depending on which type of reasoning is required. Spatial reasoning and functional reasoning require a fundamental understanding of the 3D physical world, while logical reasoning, emotional reasoning, and safety reasoning are high-level reasoning skills built upon the two fundamental reasoning abilities to address user-centric real-world applications, as shown in Fig. 2.
Spatial Reasoning measures models’ understanding of 3D spatial relationships among objects in 3D scene. It encompasses the ability to comprehend the layout of the 3D scene and the 3D location of objects within it, which could serve as the foundation for navigating or planning movements in the 3D environment.
Functional Reasoning involves understanding and inferring the purpose, function, or affordance of objects within the 3D scene. For example, functional reasoning allows an embodied agent to recognize that a chair is for sitting, a lamp is for lighting, and a refrigerator is for storing food at low temperatures. Such understanding enables the embodied agent to assist users and to perform complex tasks more effectively (e.g., turning on a lamp when the room gets dark, or navigating to a refrigerator to fetch a drink).
Logical Reasoning allows an embodied agent to not only understand its environment but also to interact with it in a goal-directed manner. For example, given a question shown in Fig. 2: “If I’m cooking dinner in the kitchen, where is the nearest place to throw the rubbish?”, an agent needs to use such reasoning ability to infer the location of objects (in this question, a rubbish bin) based on their function and spatial relationships under the specific setting (the kitchen).
Emotional Reasoning plays a critical role in human-robot interaction, where the target objects are determined by understanding human emotions, preferences, and behavioral patterns. This ability makes the embodied agents more attuned to the emotional and psychological states of humans, allowing them to provide more personalized, empathetic, and contextually appropriate responses and solutions, such as: “I’m very sad, is there something that can make me feel happy?” shown in 2.
Safety Reasoning focuses on preventing harm and ensuring the well-being of humans in the 3D environment. It requires the embodied agent to identify and assess the risk and make safety-aware decisions, such as: “Where should first aid kits be placed for easy access but out of reach of children?”.
@@ 3.3 Automatic Data Annotation with GPT-4
In total, our ScanReason dataset consists of 12929 complex reasoning question-answer-location pairs from 1456 scenes, which are split into 11455 training pairs and 1474 validation pairs. All 1474 validation question-answer pairs have been manually verified, including 342 spatial reasoning questions, 287 functional reasoning questions, 581 logical reasoning questions, 211 emotional questions, and 53 safety reasoning questions. We provide detailed statistics and more examples of our dataset in the Appendix.
@@ 3.4 Evaluation Metric
To evaluate the accuracy of predicted objects and their locations for a flexible number of ground-truth objects, we follow the evaluation metric of 3D object detection tasks. Specifically, we adopt Acc@$k$IoU as our metric, where $k$ is the threshold for the Intersection of Union (IoU) between positive predictions and ground truths. We evaluate the performance under $k=0.25$ in our experiments.
@@ 4 Method
Solving the task of 3D reasoning grounding requires the synergization of the perception, reasoning, and grounding capability of the embodied agent. Intuitively, we can first conduct reasoning based on implicit instruction such as “where should first aid kits be placed?” and the visual environment. The reasoning process provides us with information about the rough location and semantics of the object we are looking for. Then keeping that information in mind, we look back to the 3D environment to precisely locate the object. For complex scenarios, alternate reasoning and looking back are required for multiple rounds until we obtain the final answer.
Inspired by this intuition, we propose ReGround3D, consisting of a visual-centric reasoning module and a 3D grounding module with geometry-enhanced look-back, as illustrated in Fig. 3. The visual-centric reasoning module (Sec. 4.1) performs joint reasoning of language instruction and visual scene, and predicts a special <LOC> token representing the grounding information. The 3D grounding module (Sec. 4.2) looks back to the original 3D scene with comprehensive geometry information and fine-grained details. It takes the hidden embedding of the <LOC> token containing grounding-related information from the 3D features and eventually predicts the 3D locations of the target objects. Furthermore, we propose a Chain-of-Grounding mechanism (CoG) (Sec. 4.3), i.e., a chain of interleaved reasoning and grounding steps, to further synergize the grounding and reasoning capability for the 3D reasoning grounding task, as illustrated in Fig. 4.
@@ 4.1 Visual-Centric Reasoning
Specifically, for simplicity, we leverage 3D-LLM to serve as the visual-centric reasoning module because of its strong reasoning abilities inherited from the LLM. Based on the BLIP2 architecture [26], 3D-LLM uses pre-trained image encoders to extract multi-view 2D image features and back-project them into 3D spaces. The visual features are encoded by the Q-Former to produce 32 tokens as the visual input to the LLM. By leveraging the pre-trained image encoders and the Q-Former, the visual tokens encode rich semantics but lack the 3D structures, spatial interactions, and fine-grained details. Therefore, instead of directly predicting the object locations by the 3D-LLM, we ask the 3D-LLM to predict the feature representation as the output of the reasoning process, and the predicted feature is further used to ground the target object in the grounding module.
@@
3D Visual Encoder Unlike the 2D image encoder used for 3D-LLM, the 3D visual encoder directly extracts features from 3D point clouds to capture more geometric and spatial information about the 3D structures and fine-grained details. The powerful 3D visual encoder which captures comprehensive geometry, structure, layout, and fine-grained information is critical to accurate 3D grounding. Subsequently, the 3D features $f_{scene}$ produced by the 3D visual encoder and the grounding feature $h_{loc}$ from 3D-LLM are sent to the query selection module.
Query Selection Module We adopt a cross-attention mechanism, where we treat $f_{scene}$ as $Q$ (Query), and $h_{loc}$ as both $K$ (Key) and $V$ (Value), to implicitly obtain a feature-level reasoning activation heatmap. During this scene look-back process, this module roughly locates scene features that have a high response to the <LOC> token. We then select the $k$ most relevant features as the object query $f_{query}$.
3D Box Decoder is a classical transformer decoder, which consists of M transformer decoder layers, In each decoder layer, the object queries $f_{query}$ go through the text feature $h_{loc}$ cross-attention layer and scene feature $f_{scene}$ cross-attention layer. Finally, the prediction head takes the updated object queries as input and predicts the final 3D locations and matching score.
Discussion In comparison to previous works [11, 15, 21, 13] that directly predicts the bounding boxes by the LLM, the extra grounding module has the following advantages: 1) Based on the MLLM reasoning results, it allows the grounding module to perceive the scene again and focus on the region under the implicit guidance of MLLM, adapted to the user queries and the reasoning results. 2) The scene representation perceived in the grounding module can be more precise and fine-grained, which is complementary to the visual-centric reasoning module. 3) The two-step reasoning-grounding pipeline is flexible and can generalized to other types of predicting formats such as segmentation masks (by simply replacing the 3D box decoder with a 3D mask decoder).